OpenStudy (anonymous):

if u =6Bcispi/3 and v =2Bcispi/4, write u/v in polar form

OpenStudy (dumbcow):

is B a constant

OpenStudy (anonymous):

yeah

OpenStudy (dumbcow):

\[\frac{u}{v} = \frac{6B(\cos (\pi/3) +i \sin(\pi/3))}{2B(\cos (\pi/4) +i \sin(\pi/4))}\] i believe the rule when dividing vectors is subtract the angles \[\frac{u}{v} = 3[\cos(\pi/3 - \pi/4) +i \sin(\pi/3 - \pi/4)]\]

OpenStudy (anonymous):

u/v = (6B/2B)?

OpenStudy (dumbcow):

yeah you divide the scalar "r" and subtract the angles

OpenStudy (anonymous):

Oh sorry, it's meant to be 'cis' not 'cos'

OpenStudy (dumbcow):

"cis" is just shorthand for "cos + i sin"

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

didnt know that lol

OpenStudy (dumbcow):

haha that might have been something your teacher should have gone over

OpenStudy (anonymous):

so u/v = (6B/2B), cis (pi/3-pi/4) = 3cis pi/12?

OpenStudy (dumbcow):

correct

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

for the help

OpenStudy (dumbcow):

yw